OpenStudy (anonymous):

what is 2/ 45 expressed in simplest radical form?

OpenStudy (anonymous):

I'm guessing you mean \[2\sqrt{45}\] break the 45 down into its prime factors so the expression looks like this: \[2\sqrt{3*3*5}\]

OpenStudy (anonymous):

From this, we can see that 3 is under the radical twice, (as 3^2) so it can be extracted to the front as thus: \[2*3\sqrt{5}\] and you can simplify that to \[6\sqrt{5}\]

OpenStudy (anonymous):

thnxzs

OpenStudy (anonymous):

No problem :)